Grade 11Mechanics

Three stones A, B and C are simultaneously projected from same point with same speed. A is thrown upwards , B is thrown horizontally and C is thrown downwards from a building. When the distance between stone A and C becomes 10 m , then the distance between A and B will be?

PaVini GaRg
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
For A stone,
 
x1 = ut – 1/2gt^2 …..................(1)
x2 = ut + 1/2gt^2 …..................(2)
x1 + x2 = 10 …..............(3)
solving eqns. (1) (2) and (3)
2ut = 10
t = 5/u …....................................................(4)
 
For stone B,
x3 = ut
from eqn. (4)
x3 = u*5/u = 5
x3 = 5m
 
x1 = u*5/u -(1/2)*g*(25/u^2)
    = 5 – 25g/2u^2
x2 = 5 + 25g/2u^2 
 
Thus distnace between A and B will be:
root(x1^2 + x3^2)
putbthe values and get the answer.